OpenStudy (anonymous):

Simplify: 2cos^2B-cos2B

OpenStudy (ybarrap):

cos(2B) = cos^2B - sin^2B So, 2cos^2B-cos2B = 2cos^2B - (cos^2B - sin^2B) = cos^2B+sin^2B = 1

OpenStudy (anonymous):

thank you

OpenStudy (ybarrap):

You're welcome